If mass, speed and radius of rotation of a body moving on a circular path are increased by 50% then to keep the body moving in circular path increase in force required will be-

Grade:11

2 Answers

Vikas TU
7 years ago
Initially Force required that is,
F = mv^2/r
After all gets increased by 50% that is m to (m+m/2), v by (v+v/2) and r by (r + r/2) then,
new Force becomes,
F’ = (m/2)*(v^2/4)/(r/2) => mv^2/4r
dividing both we get,
F/F’ = 4
or
F’ = F/4 Hence the new fofrce must be decreased by one – fourth of the initial one.
Pawan Anil kapade
5 years ago
Initial force required,F¹=m1v1/(R1)^2m2=m1+50%m1=3/2m1v2=v1+50%v1=3/2v1r2=r1+50%r1= 3/2r1F¹/F²=4/9%change in force=[(F²-F¹)/F¹]×100%=125%